OpenStudy (anonymous):

Use an appropriate method to evalualte the limit, if it exists, of each of the following sequences {ai}. If a sequence is divergent, explain why.

OpenStudy (anonymous):

\[\frac{ 5^i-2^{3i+1} }{ 7^{i-2}+8^i }\]

OpenStudy (anonymous):

Try writing this sequence as a combination of other geometric sequences.

OpenStudy (anonymous):

What do you mean? Like manipulating it?

OpenStudy (anonymous):

Yes. Notice how each term in the numerator and denominator is some base to a power of the index? First thing you can do is break up the sequence: \[\frac{5^i-2^{3i+1}}{7^{i-2}+8^i}=\frac{5^i}{7^{i-2}+8^i}-\frac{2^{3i+1}}{7^{i-2}+8^i}\] The next thing you could do is compare to another sequence that you know converges/diverges. For example, you know that \(\dfrac{1}{8^i}\) is a convergent geometric sequence, right? Well, you know that \(\dfrac{1}{7^i+8^i}<\dfrac{1}{8^i}\), which means the left sequence also converges.

OpenStudy (anonymous):

I don't understand how you know \[\frac{ 1 }{ 7^{i}+8^i }<\frac{ 1 }{ 8^i }\]

OpenStudy (anonymous):

wait never mind okay lol .

OpenStudy (anonymous):

So for this given sequence, you know that \[\frac{5^i}{7^{i-2}+8^i}<\frac{5^i}{8^i}\] converges. Finding the limit is the tricky part, but shouldn't be too hard if you know L'Hopital's rule.

OpenStudy (anonymous):

Actually, disregard that thing I mentioned about L'Hopital's rule. You know that geometric sequences converge to 0 if it's a convergent sequence.

OpenStudy (anonymous):

So it is just 0

OpenStudy (anonymous):

Yes, it looks that way.

OpenStudy (anonymous):

At least for the first "sub-sequence" we've written. We haven't considered the one with \(2^{3i+1}\) yet.

OpenStudy (anonymous):

the top exponent is throwing me off, how would you figure out that?

OpenStudy (anonymous):

Hmm, maybe \[-\frac{2^{3i+1}}{7^{i-2}+8^i}=-2\left(\frac{2^{3i}}{\dfrac{7^i}{49}+8^i}\right)=-2\left(\frac{8^{i}}{\dfrac{7^i}{49}+8^i}\right)\] Then you dom some factoring: \[-2\left(\frac{1}{\dfrac{7^i}{49(8^i)}+1}\right)=-2\color{red}{\left(\frac{49}{\dfrac{7^i}{8^i}+49}\right)}\] The \(\dfrac{7^i}{8^i}\) approaches 0, so the red term becomes 1.

OpenStudy (anonymous):

Which means that this entire sequence approaches -2.

OpenStudy (anonymous):

Where did the 7/49 come from? D: